OpenStudy (anonymous):

Which of the following ordered pairs is a solution to the inequality y 3x - 2? (2, 0) (0, 4) (0, -3) (1, -1)

OpenStudy (anonymous):

I mean y is greater than or equal to 3x- 2

OpenStudy (anonymous):

\[(0,4)\]

OpenStudy (anonymous):

y >= 3x - 2 start subbing in with your answer choices... (2,0) 0 >= 3(2) - 2 0 >= 6 - 2 0 >= 4 (incorrect) (0,4) 4 >= 3(0) - 2 4 >= 0 - 2 4 >= - 2 (correct) (0,-3) -3 >= 3(-3) - 2 -3 >= -9 - 2 -3 >= - 11 (incorrect) (1,-1) -1 >= 3(1) - 2 -1 >= 3 - 2 -1 >= 1 (incorrect) answer : (0,4)